OpenStudy (odie8136):

Find the x- and y-intercept of the line. -7/5x -4y = 7 A. x -intercept is -7/4; y-intercept is 5 B. x -intercept is 5; y-intercept is 7/4 C. x -intercept is 5; y-intercept is -7/4 D. x -intercept is -5; y-intercept is -7/4

OpenStudy (jiteshmeghwal9):

Equation format of a line \[\frac{x}{a}+\frac{y}{b}=1\] where a is x- intercept and b is y- intercept

OpenStudy (jiteshmeghwal9):

\[\frac{-7}{5}x-4y=7\] convert it into the form \[\frac{x}{a} + \frac{y}{b} =1\]

OpenStudy (agent0smith):

To find x-int: Plug in y=0, solve for x. To find y-int: Plug in x=0, solve for y.

OpenStudy (jiteshmeghwal9):

I used the long method :p

OpenStudy (agent0smith):

Yours is way, way more complicated than necessary :P I've never even seen it.

OpenStudy (shadowlegendx):

^

OpenStudy (loveneenah):

It is X=-5(7+4y/7 and Y is Y=- 7(1+x/5)/4 Hope I helped !!